OpenStudy (anonymous):

The table shows average basketball scores per game of teams who spent a certain number of hours with each other outside of practice: Hours spent together 2 4 6 8 10 12 14 16 18 20 Average Score per game 34 39 41 41 59 58 60 68 68 80 Calculate the correlation coefficient. 1 0.97 0.87 0.77

OpenStudy (nirmalnema):

0.97

OpenStudy (nirmalnema):

Result Details & Calculation X Values ∑ = 110 Mean = 11 ∑(X - Mx)2 = SSx = 330 Y Values ∑ = 548 Mean = 54.8 ∑(Y - My)2 = SSy = 2101.6 X and Y Combined N = 10 ∑(X - Mx)(Y - My) = 808 R Calculation r = ∑((X - My)(Y - Mx)) / √((SSx)(SSy)) r = 808 / √((330)(2101.6)) = 0.9702 Meta Numerics (cross-check) r = 0.9702

OpenStudy (anonymous):

The correlation coefficent equals \[\frac{ Sxy }{ \sqrt{SxxSyy} }\] It's easier for you to google if you don't know what Sxy Sxx and Syy are if you don't know

OpenStudy (anonymous):

Thanks guys

OpenStudy (nirmalnema):

your welcome

OpenStudy (anonymous):

Correct Answer: 0.97. Took the test & got it right (: